OpenStudy (rock_mit182):

Silver dissolves in molten lead. Compute the osmotic pressure of a 0.010 m solution of silver in lead at 423°C. Compute the height of a column of molten lead (D=11.4 g cm -3) to which this pressure corresponds.

OpenStudy (rock_mit182):

@70vebelle could you help me out ?

OpenStudy (anonymous):

first use the osmotic pressure formula : We calculate the osmotic pressure, The symbol for pi (pi), using the following equation: Pi = MRT M is the molar concentration of dissolved species. R is the ideal gas constant (0.08206 L atm mol-1 K-1, or other values depending on the pressure units). T is the temperature on the Kelvin scale. Pi = 0.010M * (0.082057L-atm/mol-K) * (423+273)K Pi = 0.571 atm To find hight we have to use Pressure = height*density*gravity Height = pressure/ (desity*gravity) Here you need to convert units into SI system Convert .571 atm to SI unit for pressure P = 0.571atm*101325Pa/atm= 57868 Pascals Convert SI unit for density d = 11.4g/cm3*(1kg/1000g)(100cm/1m)^3 = 11400 kg/m3 g is in SI unit = 9.81kgm/s^2 Now you are ready to find height h = 57858 Pa /(11400 kg/m3 * 9.81kgm/s2) = 0.518 m of lead.

OpenStudy (rock_mit182):

Thanks a lot bro !